# Conservation-informed Graph Learning for Spatiotemporal Dynamics Prediction

### v1

@@ Abstract.
Data-centric methods have shown great potential in understanding and predicting spatiotemporal dynamics, enabling better design and control of the object system. However, pure deep learning models often lack interpretability, fail to obey intrinsic physics, and struggle to cope with the various domains. While geometry-based methods, e.g., graph neural networks (GNNs), have been proposed to further tackle these challenges, they still need to find the implicit physical laws from large datasets and rely excessively on rich labeled data. In this paper, we herein introduce the conservation-informed GNN (CiGNN), an end-to-end explainable learning framework, to learn spatiotemporal dynamics based on limited training data. The network is designed to conform to the general conservation law via symmetry, where conservative and non-conservative information passes over a multiscale space enhanced by a latent temporal marching strategy. The efficacy of our model has been verified in various spatiotemporal systems based on synthetic and real-world datasets, showing superiority over baseline models. Results demonstrate that CiGNN exhibits remarkable accuracy and generalization ability, and is readily applicable to learning for prediction of various spatiotemporal dynamics in a spatial domain with complex geometry.
Graph learning, spatiotemporal dynamics prediction, conservation law
@@ 1. Introduction
The principled and accurate modeling and simulation of spatiotemporal dynamical systems play an important role in many science and engineering applications, e.g., physics, meteorology, ecology, social science, material science, etc. Classical approaches are primarily rooted in the use of well-posed physics-based models, governed by a system of partial differential equations (PDEs) under certain initial conditions (ICs) and boundary conditions (BCs), where the solutions could be achieved by a properly chosen numerical solver (Ames, 2014). Nevertheless, the complexity of real-world problems poses grand challenges that traditional physics-based techniques struggle to address, especially when the prior knowledge of physics is incomplete or repeated forward analyses are required to assimilate data.
The ever-growing availability of data opens up a new avenue to tackle these challenges, resulting in data-centric methods leveraging the power of machine learning. Tremendous efforts have been recently placed on developing statistical or deep learning approaches for modeling (Brunton et al., 2020; Kochkov et al., 2021; Pan et al., 2023), predicting (Regazzoni et al., 2019; Ravuri et al., 2021; Bi et al., 2023; Zhang et al., 2023), discovering (Schmidt and Lipson, 2009; Brunton et al., 2016; Floryan and Graham, 2022), and controlling (Baggio et al., 2021; Zhai et al., 2023) the complex behavior of nonlinear dynamics. Moreover, graph neural networks (GNNs) (Wu et al., 2020; Zhang et al., 2020), which integrate topological structures with an inherent connection to the physical world, are uniquely positioned to address above issues. Above all methods learn representations from, but highly rely on, rich labeled datasets, which are often expensive to acquire from experiments or simulations in most scientific and engineering problems. Due to the model’s over-parameterized black-box nature, issues arise in the context of explainability, over-fitting, and generalizability. Embedding the prior knowledge to form physics-informed learning has shown the potential to alleviate these fundamental issues (Karniadakis et al., 2021), producing more robust models for reliable prediction (Alberts and Bilionis, 2023).
@@ 2. Related Works
Figure 1. Schematic of the CiGNN model. a, The main network architecture. b, The space block that defines a multi-mesh graph to account for different scales and transforms the physical system state to a low-dimensional graph representation. c, The SymMPNN block that is designed for learning high-dimensional graph representation. d, The Flux update component that passes symmetric and asymmetric information in graph edges. e, The time block that marches the sequenced high-dimensional features.
@@
Over the past few decades, classical numerical methods (Eymard et al., 2000; Hirani, 2003) were dominated in the forward and inverse physics problems. However, the development of this field reaches a bottleneck until the emergence of deep learning methods, which changes the status quo.
@@
As one of the most popular framework, the seminal development of the physics-informed neural network (PINN) (Raissi et al., 2019; Jagtap et al., 2020) has enabled learning in small data regimes for forward and inverse analyses of PDEs, where the loss function is regularized by the PDE residual. Such a differentiable paradigm has kindled enthusiastic attention in the past few years, which, together with its variant schemes, has been demonstrated effective in modeling and discovering a wide range of complex systems including fluid flows (Raissi et al., 2020), subsurface transport (He and Tartakovsky, 2021), cardiovascular systems (Arzani et al., 2021), deformation of solid continua (Niu et al., 2023), material constitutive relations (Rezaei et al., 2024), governing equation discovery (Chen et al., 2021), among many others. However, PINN generally has the need of explicit governing equations and the issue of scalability, resorting to the use of fully-connected neural networks and mesh-free point-wise training.
@@
Another framework, neural operator, is designed to learn the underlying Green’s functions of PDEs, offering an alternative to model spatiotemporal dynamics, which possess generalizability and do not require a prior knowledge of the explicit PDE formulation (Wu et al., 2022, 2024b). Neural operators, including DeepONet (Lu et al., 2021a), Fourier neural operator (FNO) (Li et al., 2021), and their variant algorithms (Tran et al., 2023; Lanthaler et al., 2023), have taken remarkable leaps in approximating the solution of PDE systems. However, the generalizability of these models largely depends on the richness of the labeled training data. When the complexity of the considered system remains high (e.g., multi-dimensional spatiotemporal dynamics in an irregular domain geometry), these methods might suffer from scalability issues with deteriorated performance.
@@
Recently, geometric learning of system dynamics has drawn great attention (Pineda et al., 2023; Hao et al., 2023; Wang et al., 2024), where graphs, regardless of Euclidean or non-Euclidean meshes, are employed to represent physical features and provide naturally structured relationships (e.g., spatial proximity) with interpretability (Bronstein et al., 2021; Liu et al., 2024; Wu et al., 2024a). For example, message-passing neural networks (MPNNs) have been regarded as an effective form of graph neural network (GNN) in simulating dynamics of particle assembly (Sanchez-Gonzalez et al., 2020), continuum media (Pfaff et al., 2021) and rigid bodies (Han et al., 2022), as well as forecasting global weather (Lam et al., 2023). In particular, when the explicit PDE formulation is completely unknown or partially missing, a recent study (Hernández et al., 2023) shows that forcibly encoding the law of thermodynamics into the GNN architecture can boost the generalizability of the model for simulation of quasi-static deformable bodies. Moreover, the model in (Grega et al., 2024) is designed to deal with lattice meshes and the graph solver in (Horie and Mitsume, 2024) is rooted in the assumption of conservative systems to update the discretized conservation equation in the latent space via hard encoding.
@@ 3. Preliminary and Problem Statement
Given on the theory of graph calculus, the general conservation law in Eq. 1 can be reformulated into a discrete form on graph (see Appendix Subsection B.2 for detailed derivation), shown as follows.
@@ 4. Methodology
We develop a learning scheme that conforms to the conservation law shown in Eq. 1. More details are described in Appendix Section B. We will release the source code and data at https://www.github.com after peer review.
@@ 4.1. Conservation-informed graph neural network
The CiGNN architecture is composed of five key blocks, as shown in Figure 1a, including the space block, the encoder, the SymMPNN block, the time block, and the decoder. Here, the encoder lifts the low-dimensional quantities to a high-dimensional representation, and in the meantime the decoder projects the updated high-dimensional latent features to the evolved physical state. The space block (see Figure 1b) defines a multi-mesh graph to account for different scales and transforms the physical system state to a low-dimensional graph representation (e.g., node features, graph topology). Given the fact that the disturbance of existing source, such as internal reaction, damping or external input, breaks down the conservation, the system will be subjected to entropy inequality. We introduce a multi-layer SymMPNN module (see Figure 1c–d) as the core component for high-dimensional feature representation learning, where we construct symmetric and asymmetric matrices to pass non-conservative information in graph edges. In the next work, instead of treating the SymMPNN module as a step-by-step learning and acquisition of higher-order information in space, we view the multi-layer message-passing mechanism of SymMPNN as a process of cascaded transmission over time. This implies that implicit latent time intervals, either fixed or non-fixed, can be generated between the SymMPNN layers. Hence, we propose a simple but effective latent time integrator (see the time block depicted in Figure 1e) to march the sequenced high-dimensional features, which improves the stability and accuracy of long-range multi-step rollout prediction. Moreover, we provide a pseudo-code describing our proposed graph operator, as shown in Algorithm 1. More details of design motivation are shown in Appendix Subsection B.3.
@@ 4.2.2. Processor
where $\mathbf{h}_{i}^{l}\in\mathbb{R}^{c}$ and $\mathbf{h}_{j}^{l}\in\mathbb{R}^{c}$ denote the high-dimensional node features at nodes $i$ and $j$, respectively, at the $l$th SymMPNN layer; $\mathbf{e}_{ij}^{l}\in\mathbb{R}^{c}$ the high-dimensional edge features between the $i$th and $j$th nodes; $\mathbf{e}^{l,*}_{ij}\in\mathbb{R}^{c}$ and $\mathbf{e}^{l,*}_{ji}\in\mathbb{R}^{c}$ the corresponding intermediate edge features (e.g., bidirectional) inferred to represent symmetric and asymmetric information passing; $\phi_{e}^{l}(\cdot)$ the edge update function (e.g., MLP); $\mathbf{w}^{l}\in\mathbb{R}^{c}$ the trainable weights shared over the entire graph for the $l$th layer; $\mathcal{N}_{i}$ the neighborhood node’s index set for node $i$; and $\circ$ the Hadamard product. Here, $\mathbf{e}^{l+1}_{ij}$ facilitates the learning process by strictly obeying the conservation law and the inequality principle. $\phi_{v}^{l}(\cdot)$ is a node update function (e.g., MLP) and $\phi^{l}(\cdot)$ a flux update function (e.g., MLP) for symmetric message passing.
We found that, when $\mathbf{w}^{l}\equiv 1$, the proposed CiGNN model achieves a more reliable performance. For comparison purpose, we remove the symmetric update part as shown in Eq. 5b while only keeping the asymmetric portion. Thus, we obtain two variant models, namely, CiGNN- ($\mathbf{w}^{l}\equiv 1$) and CiGNN ($\mathbf{w}^{l}\neq 1$, which is trainable).
@@ 4.3.1. Space Block with multi-mesh strategy
Consider a graph with $G=(V,E)$, where $V$ denotes the node set and $E$ the edge set. The multi-mesh strategy involves three steps: (1) defining $G^{k}$ as the initial graph; (2) randomly discarding a proportion of nodes in $G^{k}$ to generate a new graph $G^{k+1}$ as the subsequent initial graph; (3) iteratively repeating the second step, and aggregating all graphs into $G$. We organize the generated graphs by layers, designating the graph with the most node count as the topmost layer and the one with the fewest nodes as the bottom layer. In other words, as $k$ increases ($k=0,1,2,...$), the resulting graph structure $G^{k+1}$ contains fewer nodes. Then, we define a refinement ratio $r\in[0,1]$ to characterize the relationship between node numbers in the context graph. For instance, if $G^{k}$ has $N$ nodes, then the subsequent graph, $G^{k+1}$, will have $rN$ nodes. We specifically set the refinement ratio $r$ to 0.1 in this study. Experimental results concerning the selection of $r$ are discussed in Appendix Table S7. The overall graph $G$ can be conceptualized as the union of all edges and nodes:
@@
Table 1. Summary of each model’s performance in terms of prediction error, e.g., aggregated root mean square error (RMSE) between the predicted and ground truth test data. Here, “–” denotes the model is unable or unsuitable to learn the dynamics (e.g., inapplicable to handle irregular mesh). “MP-PDE” means that no gradient is backward propagated except for the last step. The bold values and underlined values represent the optimal and sub-optimal results on various datasets. “Pro.” means the promotion value which is calculated from the optimal and sub-optimal results. The one-step prediction results are shown in Appendix Table S9.
@@ 4.3.2. Time Block with trainable temporal integration
Notably, this temporal integration strategy is realized on the high-dimensional space, so the coefficients $\alpha_{j},\beta_{j}$ are also learnable latent features. $\mathbf{K}_{L}$ denotes the output of final SymMPNN layer, which is the $L$-order approximation of solution. Each increment in our method can be viewed as the product of the coefficients $\beta_{j}$ and an estimated slope specified by each function $\varphi_{j}$. The total increment is the sum of the weighted increments. To this end, we employ a trainable 1D convolutional filter to approximate the coefficients in Eqs. 8a and 8b, as shown in Figure 1d. This temporal integration strategy does not lead to extra cost for network training and inference.
Figure 2. Diagrams of irregular domains. a, Schematic diagram of the CF example setup with the fluid flowing in from the left (inflow) to the right (outflow). b, The geographic and sensor information of the BS Real-world dataset.
@@ 5. Experiments
To verify the efficacy of CiGNN for learning complex spatiotemporal dynamics on graphs, we consider four different systems shown in Appendix Table S2, e.g., the 2D Burgers equation, a 3D Gray-Scott (GS) reaction-diffusion (RD) system, the 2D incompressible flow past a cylinder (CF), and the Black Sea (BS) hydrological observation (see Figure 2). The datasets of the first three systems were spatiotemporally down-sampled (e.g., 5-fold in time) from high-fidelity data generated by direct numerical simulations considering different ICs, BCs, and meshes, while the last one was compiled from field measurements. Additional information, including data generation statistics and example trajectories, can be found in Appendix Table S3. We selected four baseline models for comparison (see Appendix Subsection C.3 for more details), including deep operator network (DeepONet) (Lu et al., 2021b), Fourier neural operator (FNO) (Li et al., 2021), mesh-based graph network (MGN) (Pfaff et al., 2021), and message-passing PDE solver (MP-PDE) (Brandstetter et al., 2022), with hyperparameters listed in Appendix Tables S11S13.
5.2. Loss function
The training procedure is to minimize the discrepancy between the labeled and the predicted data. The loss function of our model is defined as
(9) $\mathcal{L}(\bm{\theta})=\frac{1}{dSN\widetilde{T}}\sum_{\alpha=1}^{S}\left\|vec\big{(}\widehat{\mathbf{U}}^{(\alpha)}\big{)}-vec\big{(}\mathbf{U}^{(\alpha)}\big{)}\right\|_{2}^{2},$
which calculates the aggregated mean square errors (MSE) between the rollout-predicted data $\widehat{\mathbf{U}}^{(\alpha)}\in\mathbb{R}^{N\times d\times\tilde{T}}$ and the training labels $\mathbf{U}^{(\alpha)}\in\mathbb{R}^{N\times d\times\tilde{T}}$ to optimize the trainable parameters ($\bm{\theta}$) in our model. Here, $vec(\cdot)$ denotes the vectorization operation; $d$ the system dimension; $N$ the total number of mesh grid points; $\widetilde{T}$ the number of rollout time steps in training; and $S$ the total number of grouped training samples.
5.3. Evaluation metrics
The root mean squared error (RMSE) and the Pearson correlation coefficient (PCC) are utilized as the evaluation metrics, given by

### v2

@@ Abstract.
Data-centric methods have shown great potential in understanding and predicting spatiotemporal dynamics, enabling better design and control of the object system. However, deep learning models often lack interpretability, fail to obey intrinsic physics, and struggle to cope with the various domains. While geometry-based methods, e.g., graph neural networks (GNNs), have been proposed to further tackle these challenges, they still need to find the implicit physical laws from large datasets and rely excessively on rich labeled data. In this paper, we herein introduce the conservation-informed GNN (CiGNN), an end-to-end explainable learning framework, to learn spatiotemporal dynamics based on limited training data. The network is designed to conform to the general conservation law via symmetry, where conservative and non-conservative information passes over a multiscale space enhanced by a latent temporal marching strategy. The efficacy of our model has been verified in various spatiotemporal systems based on synthetic and real-world datasets, showing superiority over baseline models. Results demonstrate that CiGNN exhibits remarkable accuracy and generalizability, and is readily applicable to learning for prediction of various spatiotemporal dynamics in a spatial domain with complex geometry.
Graph neural network, spatiotemporal dynamics prediction, conservation law
@@ 1. Introduction
The principled and accurate modeling and simulation of spatiotemporal dynamical systems play an important role in many science and engineering applications, e.g., physics, meteorology, ecology, social science, material science, etc. Classical approaches are primarily rooted in the use of well-posed physics-based models, governed by a system of partial differential equations (PDEs) under certain initial conditions (ICs) and boundary conditions (BCs), where the solutions could be achieved by a properly chosen numerical solver (Ames, 2014). Nevertheless, the complexity of real-world problems poses grand challenges that traditional physics-based techniques struggle to address, especially when the physical prior knowledge is incomplete or repeated forward analyses are required to assimilate data.
The ever-growing availability of data opens up a new avenue to tackle these challenges, resulting in data-centric methods leveraging the power of machine learning. Tremendous efforts have been recently placed on developing statistical or deep learning approaches for modeling (Brunton et al., 2020; Kochkov et al., 2021; Pan et al., 2023), predicting (Regazzoni et al., 2019; Ravuri et al., 2021; Bi et al., 2023; Zhang et al., 2023), discovering (Schmidt and Lipson, 2009; Brunton et al., 2016; Floryan and Graham, 2022), and controlling (Baggio et al., 2021; Zhai et al., 2023) the complex behavior of nonlinear dynamics. Moreover, graph neural networks (GNNs) (Wu et al., 2020; Zhang et al., 2020a), which integrate topological structures with an inherent connection to the physical world, are uniquely positioned to address above issues. Above all methods learn representations from, but highly rely on, rich labeled datasets, which are often expensive to acquire from experiments or simulations in most scientific and engineering problems. Due to the model’s over-parameterized black-box nature, issues arise in the context of explainability, over-fitting, and generalizability. Embedding the prior knowledge to form physics-informed learning has shown the potential to alleviate these fundamental issues (Karniadakis et al., 2021), producing more robust models for reliable prediction (Alberts and Bilionis, 2023).
@@ 2. Related Works
Figure 1. Schematic of the CiGNN model. a, The main network architecture. b, The space block that defines a multi-mesh graph to account for different scales and transforms the physical state to a low-dimensional representation. c, The SymMPNN block that is designed for learning high-dimensional representation. d, The Flux update component that passes symmetric and asymmetric information on edges. e, The time block that marches the sequenced high-dimensional features.
@@
Over the past few decades, classical numerical methods (Eymard et al., 2000; Hirani, 2003) were dominated in the forward and inverse physics problems. However, the development of this field reaches a bottleneck until the deep learning methods’ emergence, which changes the status quo.
@@
As one of the most popular framework, the seminal development of the physics-informed neural network (PINN) (Raissi et al., 2019; Jagtap et al., 2020) has enabled learning in small data regimes for forward and inverse analyses of PDEs, where the loss function is regularized by the PDE residual. Such a differentiable paradigm has kindled enthusiastic attention in the past few years, which, together with its variant schemes, has been demonstrated effective in modeling and discovering a wide range of complex systems including fluid flows (Raissi et al., 2020), subsurface transport (He and Tartakovsky, 2021), cardiovascular systems (Arzani et al., 2021), solid continua’s deformation (Niu et al., 2023), material constitutive relations (Rezaei et al., 2024), governing equation discovery (Chen et al., 2021), data superresolution (Ren et al., 2023), elastic waves (Rao et al., 2021), structural metamodeling (Zhang et al., 2020b, c), among many others. However, PINN generally has the need of explicit equations and the issue of scalability, resorting to the use of fully-connected neural networks and mesh-free point-wise training.
@@
Another framework, neural operator, is designed to learn the underlying Green’s functions of PDEs, offering an alternative to model spatiotemporal dynamics, which possess generalizability and do not require a prior knowledge of the explicit PDE formulation (Wu et al., 2022, 2024b). Neural operators, including DeepONet (Lu et al., 2021a), Fourier neural operator (FNO) (Li et al., 2021), and their variant algorithms (Tran et al., 2023; Lanthaler et al., 2023; Wang et al., 2024b), have taken remarkable leaps in approximating the solution of PDE systems. However, the generalizability of these models largely depends on the richness of the labeled training data. When the complexity of the considered system remains high (e.g., multi-dimensional spatiotemporal dynamics in an irregular domain geometry), these methods might suffer from scalability issues with deteriorated performance.
@@
Recently, geometric learning of system dynamics has drawn great attention (Pineda et al., 2023; Hao et al., 2023; Wang et al., 2024a), where graphs, regardless of Euclidean or non-Euclidean meshes, are employed to represent physical features and provide naturally structured relationships (e.g., spatial proximity) with interpretability (Liu et al., 2024; Wu et al., 2024a). For example, message-passing neural networks (MPNNs) have been regarded as an effective form of graph neural network (GNN) in simulating dynamics of particle assembly (Sanchez-Gonzalez et al., 2020), continuum media (Pfaff et al., 2021) and rigid bodies (Han et al., 2022), as well as forecasting global weather (Lam et al., 2023). In particular, when the explicit PDE formulation is completely unknown or partially missing, a recent study (Hernández et al., 2023) shows that forcibly encoding the law of thermodynamics into the GNN architecture can boost the generalizability of the model for simulation of quasi-static deformable bodies. Moreover, the model in (Grega et al., 2024) is designed to deal with lattice meshes and the graph solver in (Horie and Mitsume, 2024) is rooted in the assumption of conservative systems to update the discretized conservation equation in the latent space via hard encoding.
@@ 3. Preliminary and Problem Statement
Given on the theory of graph calculus, the general conservation law in Eq. 1 can be reformulated into a discrete form on graph (see Appendix Section A for detailed derivation), shown as follows.
@@ 4. Methodology
We develop a learning scheme that conforms to the conservation law shown in Eq. 1. The schematic overview of our proposed CiGNN is shown in Figure 1 which illustrates an end-to-end learning pipeline of “Encoder-Processor-Decoder” framework.
@@ 4.1. Conservation-informed graph neural network
The CiGNN architecture is composed of five key blocks, as shown in Figure 1a, including the space block, the encoder, the SymMPNN block, the time block, and the decoder. Here, the encoder lifts the low-dimensional quantities to a high-dimensional representation, and in the meantime the decoder projects the updated high-dimensional latent features to the evolved physical state. The space block (see Figure 1b) defines a multi-mesh graph to account for different scales and transforms the physical system state to a low-dimensional graph representation (e.g., node features, graph topology). Given the fact that the disturbance of existing source, such as internal reaction, damping or external input, breaks down the conservation, the system will be subjected to entropy inequality. We introduce a multi-layer SymMPNN module (see Figure 1c–d) as the core component for high-dimensional feature representation learning, where we construct symmetric and asymmetric matrices to pass non-conservative information in graph edges. In the next work, instead of treating the SymMPNN module as a step-by-step learning and acquisition of higher-order information in space, we view the multi-layer message-passing mechanism of SymMPNN as a process of cascaded transmission over time. This implies that implicit latent time intervals, either fixed or non-fixed, can be generated between the SymMPNN layers. Hence, we propose a simple but effective latent time integrator (see the time block depicted in Figure 1e) to march the sequenced high-dimensional features, which improves the stability and accuracy of long-range multi-step rollout prediction. Moreover, we provide a pseudo-code describing our proposed graph operator, as shown in Algorithm 1.
@@ 4.2.2. Processor
where $\mathbf{h}_{i}^{l}\in\mathbb{R}^{c}$ and $\mathbf{h}_{j}^{l}\in\mathbb{R}^{c}$ denote the high-dimensional node features at nodes $i$ and $j$, respectively, at the $l$th SymMPNN layer; $\mathbf{e}_{ij}^{l}\in\mathbb{R}^{c}$ the high-dimensional edge features between the $i$th and $j$th nodes; $\mathbf{e}^{l,*}_{ij}\in\mathbb{R}^{c}$ and $\mathbf{e}^{l,*}_{ji}\in\mathbb{R}^{c}$ the corresponding intermediate edge features (e.g., bidirectional) inferred to represent symmetric and asymmetric information passing; $\phi_{e}^{l}(\cdot)$ and $\phi_{v}^{l}(\cdot)$ the edge and node update functions (e.g., MLPs); $\mathbf{w}^{l}\in\mathbb{R}^{c}$ the trainable weights shared over the entire graph for the $l$th layer; $\mathcal{N}_{i}$ the neighborhood node’s index set for node $i$; and $\circ$ the Hadamard product. Here, $\mathbf{e}^{l+1}_{ij}$ facilitates the learning process by strictly obeying the conservation law and the inequality principle. $\phi^{l}(\cdot)$ a flux update function (e.g., MLP) for symmetric message passing.
We found that, when $\mathbf{w}^{l}\equiv 1$, our CiGNN model achieves a more reliable performance. For comparison purpose, we remove the symmetric update part as shown in Eq. 5b while only keeping the asymmetric portion. Thus, we obtain two variant models, namely, CiGNN- ($\mathbf{w}^{l}\equiv 1$) and CiGNN ($\mathbf{w}^{l}\neq 1$, which is trainable).
@@ 4.3.1. Space Block with multi-mesh strategy
Consider a graph with $G=(V,E)$, where $V$ denotes the node set and $E$ the edge set. The multi-mesh strategy involves three steps: (1) defining $G^{k}$ as the initial graph; (2) randomly discarding a proportion of nodes in $G^{k}$ to generate a new graph $G^{k+1}$ as the subsequent initial graph; (3) iteratively repeating the second step, and aggregating all graphs into $G$. We organize the generated graphs by layers, designating the graph with the most node count as the topmost layer and the one with the fewest nodes as the bottom layer. In other words, as $k$ increases ($k=0,1,2,...$), the resulting graph structure $G^{k+1}$ contains fewer nodes. Then, we define a refinement ratio $r\in[0,1]$ to characterize the relationship between node numbers in the context graph. For instance, if $G^{k}$ has $N$ nodes, then the subsequent graph, $G^{k+1}$, will have $rN$ nodes. We specifically set the refinement ratio $r$ to 0.1 in this study. The overall graph $G$ can be conceptualized as the union of all edges and nodes:
@@
Table 1. Summary of each model’s performance in terms of prediction error, e.g., aggregated root mean square error (RMSE) between the predicted and ground truth test data. Here, “–” denotes the model is unable or unsuitable to learn the dynamics (e.g., inapplicable to handle irregular mesh). “MP-PDE” means that no gradient is backward propagated except for the last step. The bold values and underlined values represent the optimal and sub-optimal results on various datasets. “Pro.” means the promotion value which is calculated from the optimal and sub-optimal results.
@@ 4.3.2. Time Block with trainable temporal integration
Notably, this temporal integration strategy is realized on the high-dimensional space, so the coefficients $\alpha_{j},\beta_{j}$ are also learnable latent features. $\mathbf{K}_{L}$ denotes the output of final SymMPNN layer, which is the $L$-order approximation of solution. Each increment can be viewed as the product of the coefficients $\beta_{j}$ and an estimated slope specified by each function $\varphi_{j}$. The total increment is the sum of the weighted increments. To this end, we employ a trainable 1D convolutional filter to approximate the coefficients in Eqs. 8a and 8b, as shown in Figure 1d. This temporal integration strategy does not lead to extra cost for network training and inference.
Figure 2. Diagrams of irregular domains. a, Schematic diagram of the CF example setup with the fluid flowing in from the inflow to the outflow. b, The geographic and sensor information of the BS Real-world dataset.
@@ 5. Experiments
Table 2. Basic information of datasets.
3D GS RD Cubic $[0,96]^{3}$ ($u,v$) 13,824 ($24^{3}$) 3,000 10 (5/3/2) Periodic No
@@
To verify the efficacy of CiGNN for learning complex spatiotemporal dynamics on graphs, we consider four different systems shown in Table 2, e.g., the 2D Burgers equation, a 3D Gray-Scott (GS) reaction-diffusion (RD) system, the 2D incompressible flow past a cylinder (CF), and the Black Sea (BS) hydrological observation (see Figure 2). The datasets of the first three systems were spatiotemporally down-sampled (e.g., 5-fold in time) from high-fidelity data generated by direct numerical simulations considering different ICs, BCs, and meshes, while the last one was compiled from field measurements. More detailed information of data generation can be found in Appendix Section B. We selected four baseline models for comparison , including deep operator network (DeepONet) (Lu et al., 2021b), Fourier neural operator (FNO) (Li et al., 2021), mesh-based graph network (MGN) (Pfaff et al., 2021), and message-passing PDE solver (MP-PDE) (Brandstetter et al., 2022).
5.2. Loss function and Evaluation metrics
The training procedure is to minimize the discrepancy between the labeled and the predicted data. The loss function is defined as
(9) $\mathcal{L}(\bm{\theta})=\frac{1}{dN\widetilde{T}}\left\|vec\big{(}\widehat{\mathbf{U}}\big{)}-vec\big{(}\mathbf{U}\big{)}\right\|_{2}^{2},$
which calculates the aggregated mean square errors (MSE) between the rollout-predicted data $\widehat{\mathbf{U}}\in\mathbb{R}^{N\times d\times{T}}$ and the training labels $\mathbf{U}\in\mathbb{R}^{N\times d\times{T}}$ to optimize the trainable parameters ($\bm{\theta}$) in our model. Here, $vec(\cdot)$ denotes the vectorization operation; $d$ the system dimension; $N$ the total number of mesh grid points; ${T}$ the number of rollout time steps in training. The root mean squared error (RMSE) and the Pearson correlation coefficient (PCC) are utilized as the evaluation metrics, given by